OpenStudy (anonymous):

When the quadratic equation y = (x − 3)(4x + 2) is written in standard form, what is the value of the coefficient "b?"

OpenStudy (anonymous):

\[y=ax^2+bx+c\]

OpenStudy (anonymous):

that's standard form.

OpenStudy (anonymous):

you have to distribute out \[y = (x − 3)(4x + 2) \]

OpenStudy (anonymous):

4x^2 -10x -6 -10 is b ( use the standard form)

hartnn (hartnn):

simply multiply term by term,and compare with ax*x+bx+c...so 4x*x+(-12+2)x-6...so b is -10

OpenStudy (anonymous):

wow give answers much guys?

OpenStudy (amistre64):

y = (a+b)(m+n) = am + an + bm + bn = am + (an + bm) + bn

OpenStudy (farmdawgnation):

@hartnn Please try to engage the user in learning, instead of just giving them the answer in the future. Thanks.

hartnn (hartnn):

roger that

OpenStudy (lgbasallote):

sadly @shayyeatsrainbows 's not engaging either...